# Queryable Prototype Multiple Instance Learning with Vision-Language Models for Incremental Whole Slide Image Classification

### v1

@@ Conclusions
This paper proposes the first Vision-Language-based framework for incremental WSI classification, QPMIL-VL, breaking through traditional vision modality. By a pool of learnable prototype pairs, it aggregates the instance features from pre-trained image encoder to generate bag-level feature under the guidance of matched prototype prompts. Class probability is predicted by the Consine similarity between bag-level feature and enhanced class feature. Extensive experiments demonstrate that our method and its components are effective. It obviously outperforms other methods in incremental WSI classification, achieving SOTA results.

### v2

@@ Conclusions
This paper proposes the first Vision-Language-based framework for incremental WSI classification, QPMIL-VL, breaking through traditional vision modality. By a pool of learnable prototype pairs, it aggregates the instance features from pre-trained image encoder to generate bag-level feature under the guidance of matched prototype prompts. Class probability is predicted by the cosine similarity between bag-level feature and enhanced class feature. Extensive experiments demonstrate that our method and its components are effective. It obviously outperforms other methods in incremental WSI classification, achieving SOTA results.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No.62476049.